OpenStudy (anonymous):

Let z=2-5i a)Write z in polar form. b) Find the complex number z^3 and express it in polar form. c) Find the three cube roots of z and express them in polar form.

OpenStudy (anonymous):

When \(z=a+bi\), \[r=\sqrt{a^2+b^2}\]and\[\theta_r=tan\frac{b}{a}\]But let's draw it!We know it's in the lower-right quadrant, so our \(\theta\) will be given by \(2\pi - \theta_r\). Finally, what does the final form look like? \[z=r(cos\theta + isin\theta)\]